# On the nullspace of split graphs

### v2

@@ 4 Unbalanced split graphs
$G_{v*}x=x_{u}+\sum_{i\in N(v)\setminus\{u\}}x_{i}=0=G_{u*}x=x_{v}+\sum_{i\in N(u)\setminus\{v\}}x_{i}.$
Since $N(v)\setminus\{u\}=N(u)\setminus\{v\}$, we have
$\sum_{i\in N(v)\setminus\{u\}}x_{i}=\sum_{i\in N(u)\setminus\{v\}}x_{i}.$
@@
Let $Sp(K,S)$ be an unbalanced split graph. Then, we have the following:
@@
    It follows from Theorem 4.4, Theorem 4.2(2) and Theorem 4.1(1).
@@
    From Theorem 4.2(2) we know that $\mathcal{S}(Sp)$ contains all the maximum independent sets of $Sp$. Thus, from Theorem 4.1(2) we easily derive that the intersection of all of them is $S^{*}(Sp)$. Hence, $\operatorname{Supp}(Sp)\subset S^{*}(Sp)$ by Theorem 4.4, and so $W(Sp)\cap\operatorname{Supp}(Sp)=\varnothing$ by (2).
    Alternatively: if $x\in\operatorname{\mathcal{N}}(Sp)$, then $x_{u}=x_{v}$ for all $u,v\in W(Sp)$, by Lemma 4.7 and Lemma 4.6. Since $\operatorname{Supp}(Sp)$ is independent by Theorem 4.4, it must be $x_{w}=0$ for all $w\in W(Sp)$, because $|W(Sp)|\geq 2$ and $W(Sp)$ is a clique. Consequently, $W(Sp)\cap\operatorname{Supp}(Sp)=\varnothing$.
@@
    If $u,v$ are distinct swing vertices, observe that $A(Sp)_{*u}=A(Sp)_{*v}$. As a consequence, $e_{u}-e_{v}\in\operatorname{\mathcal{N}}(Sp)$. Then, $W(Sp)\subset\operatorname{Supp}(Sp)$.
    On the other hand, we know from Theorem 4.1(3) and Theorem 4.2(2) that $S^{*}(Sp)\cup W(Sp)$ is the unique maximum independent set of $Sp$. Therefore, $\operatorname{Supp}(Sp)\subset S^{*}(Sp)\cup W(Sp)$, by Theorem 4.4.
Proposition 4.9.
Let $Sp(K,S)$ be a connected split graph such that $K=\bigcup_{v\in S}N_{Sp}(v)$. If $\operatorname{rk}(R)=1$, then $R=J$. Moreover, all the independent (clique) vertices of $Sp$ are pairwise twins.
Proof.
Since $Sp$ is connected and has no isolated vertices in $S$, all the columns of $R$ are nonzero. Since $K=\bigcup_{v\in S}N(v)$, all the rows of $R$ are nonzero. Moreover, since $\operatorname{rk}(R)=1$, all the columns of $R$ must be scalar multiples of each other. Since all columns are nonzero $\{0,1\}$-vectors, they must all be equal to the same vector. Since $R_{i*}\neq 0$ for all $i$ and $R$ is a $\{0,1\}$-matrix, we get that $R_{*j}=\operatorname{1\!\!1}$ for all $j$. Therefore, $R=J$.
Since $R=J$, every vertex in $S$ is adjacent to all vertices in $K$, i.e., $N(v)=K$ for all $v\in S$. Therefore, all the independent vertices are pairwise twins. Similarly, every vertex in $K$ is adjacent to all vertices in $S$ and all other vertices in $K$. Hence,
$N(u)\setminus\{v\}=(K\setminus\{u,v\})\cup S=N(v)\setminus\{u\}$
for all $u,v\in K$, which means that all the clique vertices are pairwise twins. ∎
@@ 5 Split graphs with supported clique vertices
In this section we examine the structure of the nullspace when clique vertices are involved; this scenario gives rise to rich yet highly constrained linear dependencies. We prove that the subspace of kernel vectors with support in the clique—the clique-kernel —is at most one-dimensional. This foundational result leads to a precise dichotomy: the overall nullity of the split graph is either equal to the nullity of its biadjacency matrix $R$, or exactly one greater. This establishes a complete and computable classification of the nullity based purely on the dimension of this key subspace, tightly linking the combinatorial presence of supported clique vertices to the algebraic property of the graph’s nullity.
The following result is crucial as it shows that when clique vertices appear in the support, they do so in essentially a one-dimensional way, which has important implications for the structure of the nullspace.
@@
Now we can show that the appearance of clique vertices in the support is an ”all-or-nothing” phenomenon in a one-dimensional sense, which has significant implications for understanding the structure of kernel vectors in split graphs.
@@
We obtain the desired result by combining item (3) of Theorem 3.4 with Lemma 5.1. ∎
@@
By Lemma 5.1, $\dim(\operatorname{Cker}{\left(Sp\right)})\leq 1$. If $\dim(\operatorname{Cker}{\left(Sp\right)})=0$, then by Theorem 3.4(3), $\operatorname{Supp}(Sp)\subset S$, and hence $\operatorname{nl}(Sp)=\operatorname{nl}(R)$ by Theorem 3.2(3). If $\dim(\operatorname{Cker}{\left(Sp\right)})=1$, then by Theorem 5.2, $\operatorname{Supp}(Sp)\cap K\neq\varnothing$, and hence $\operatorname{nl}(Sp)=\operatorname{nl}(R)+1$ by Theorem 5.3. ∎
@@ 6 Split graphs with nullity 1
Having determined the general structure of the nullspace, we now specialize in split graphs whose nullity is exactly one. This condition imposes a remarkable rigidity on the graph’s structure, forcing a highly constrained relationship between the kernel, the adjugate of the adjacency matrix, and the graph’s support. In this section, we leverage the classical properties of the adjugate for singular matrices to derive precise combinatorial descriptions of these graphs, culminating in formulas that link the coordinates of the unique kernel vector to the determinants of principal submatrices.
@@
So, $a_{i}\in\operatorname{\mathcal{N}}(A)$ for all $i\in[n]$, from which we deduce that $\operatorname{rk}(\operatorname{adj}(A))\leq 1$, because $\operatorname{nl}(A)=1$. On the other hand, we know by Lemma 6.1 that $\operatorname{rk}(\operatorname{adj}(A))\neq 0$. Therefore, $\operatorname{rk}(\operatorname{adj}(A))=1$. ∎
@@
    Let $\operatorname{\mathcal{N}}(G)=\langle z\rangle$ and $s=|\operatorname{Supp}(G)|$. First, we label $V(G)$ using the ordering $(\operatorname{Supp}(G),V(G)\setminus\operatorname{Supp}(G))$. By this labeling, $z=(x,0)^{t}$, where $x\in\mathbb{R}^{s}$ and $x_{v}\neq 0$ for all $v\in[s]$. Then, by Lemma 6.2, we can write
@@
From Lemma 6.5, we know that $B_{i}$ spans the subspace $\langle\{A_{*j}:j\in\operatorname{Supp}(A)\}\rangle$. With this in mind, it is clear that $B$ spans $\operatorname{\mathcal{R}}(A)$. Since $|B|=n-1=\operatorname{rk}(A)$, we can finally conclude that $B$ is a basis for $\operatorname{\mathcal{R}}(A)$. ∎
@@
This theorem provides a very concrete and useful result: for any graph with nullity 1 and no isolated vertices, you can obtain a basis for the column space by simply removing any one column corresponding to a supported vertex. This is computationally useful and reveals the highly structured nature of the adjacency matrix in this case.
@@
Since $G$ has no isolated vertices, $A(G)_{*j}\neq 0$ for all $j\in V(G)$. If $\operatorname{Supp}(G)=\{u\}$ for some $u$, then $\operatorname{\mathcal{N}}(G)=\langle e_{u}\rangle$. But, $0=A(G)e_{u}=A(G)_{*u}$, contradicting that all the columns of $A(G)$ are nonzero. Therefore, $|\operatorname{Supp}(G)|\geq 2$, and so we can apply Lemma 6.6. ∎
@@
Since $P$ is obtained by selecting a basis for $\operatorname{\mathcal{R}}(R)$, we have $\operatorname{\mathcal{R}}(P)=\operatorname{\mathcal{R}}(R)$ (hence, $\operatorname{\mathcal{R}}((I-J)^{-1}P)=\operatorname{\mathcal{R}}((I-J)^{-1}R)$) and $\operatorname{\mathcal{N}}(P^{t})=\operatorname{\mathcal{N}}(R^{t})$. Therefore, $\operatorname{Cker}{\left(Sp\right)}=\operatorname{Cker}{\left(Sp^{\prime}\right)}$. Since $\operatorname{Cker}{\left(Sp\right)}\neq\{0\}$, by Theorem 3.4(3), we have $\operatorname{Supp}(Sp^{\prime})$ $\cap K\neq\varnothing$. Since $\operatorname{\mathcal{N}}(P)=\{0\}$, applying Corollary 6.4 to $Sp^{\prime}$ gives $\operatorname{nl}(Sp^{\prime})=1$. ∎
@@
    The equality $\operatorname{Supp}(Sp^{\prime})\cap K=\operatorname{Supp}(Sp)\cap K$ can easily be derived from Lemma 6.8. To obtain that $\operatorname{Supp}(Sp^{\prime})\cap S^{\prime}\subset\operatorname{Supp}(Sp)\cap S^{\prime}$, recall that, by item (2) of Theorem 3.4 and item (1) of Theorem 3.2, there exists a (unique) vector $y_{0}$ such that $(I-J)^{-1}Py_{0}=z$. Then,
@@
    Let $M$ be the $|V(Sp)|\times(|V(Sp^{\prime})|-1)$ matrix obtained from $A(Sp)$ by deleting its last $|S|-\operatorname{rk}(R)$ columns and the column $s$. Since $\operatorname{rk}(M)$ is the size of its largest invertible submatrix, we have that $\operatorname{rk}(M)=|V(Sp^{\prime})|-1$ because $M$ contains $A(Sp^{\prime}-s)$, which is nonsingular by Lemma 6.8 and Theorem 6.3. Hence, $B_{s}$ is a linearly independent set.
@@ 7 Tyshkevich composition
Understanding how the nullspace behaves under graph operations is a natural and important question. In this section, we analyze how the nullspace of a split graph interacts with the Tyshkevich composition, a well-known product that builds a larger graph by fully joining the clique of a split graph to an arbitrary second graph. We establish fundamental results showing how the singularity of the composition is controlled by the properties of its components, providing a powerful tool for constructing graphs with prescribed nullity and for understanding the propagation of kernel vectors through this structured gluing.
@@
This operation is particularly relevant for studying the nullspace of split graphs, as it allows us to build more complex split graphs from simpler components while potentially controlling spectral properties.
@@ 8 Determinant of split graphs
Thanks to the block structure of their adjacency matrices, split graphs admit explicit combinatorial determinant formulas that can be obtained with the Schur complement and the matrix-determinant lemma.
@@
Theorem 8.2 (Matrix determinant lemma).
If $M$ is an invertible matrix and $u,v$ are vectors, then
$\det(M+uv^{t})=(1+v^{t}M^{-1}u)\det(M).$
@@
$\det(Sp)=(-1)^{k-1}(k-1)\det\left(R^{t}R-\frac{rr^{t}}{k-1}\right).$
Furthermore, if $\operatorname{\mathcal{N}}(R)=\{0\}$, we have
$\det\left(R^{t}R-\frac{rr^{t}}{k-1}\right)=\left(1-\frac{1}{k-1}r^{t}(R^{t}R)^{-1}r\right)\det(R^{t}R).$
@@
Using the fact that $(I-J)^{-1}=I-(k-1)^{-1}J$, we have:
@@
For the second part, if $\operatorname{\mathcal{N}}(R)=\{0\}$, then $R^{t}R$ is invertible, and we apply the matrix determinant lemma (Theorem 8.2) with $M=R^{t}R,u=\frac{r}{1-k}$ and $v=r$.
$\det\left(R^{t}R-\frac{rr^{t}}{k-1}\right)=\left(1-\frac{1}{k-1}r^{t}(R^{t}R)^{-1}r\right)\det(R^{t}R).\qed$
@@
Let $Sp(K,S)$ be a split graph such that $|K|\geq 2$ and $\operatorname{\mathcal{N}}(R)=\{0\}$. Then, $Sp$ is singular if and only if
$\operatorname{1\!\!1}^{t}R(R^{t}R)^{-1}R^{t}\operatorname{1\!\!1}=|K|-1.$
@@
It immediately follows from Theorem 8.3, because $\det(R^{t}R)\neq 0$. ∎

### v3

@@ 4 Unbalanced split graphs
$G_{v*}x=x_{u}+\sum_{i\in N(v)\setminus u}x_{i}=0=G_{u*}x=x_{v}+\sum_{i\in N(u)\setminus v}x_{i}.$
Since $N(v)\setminus u=N(u)\setminus v$, we have
$\sum_{i\in N(v)\setminus u}x_{i}=\sum_{i\in N(u)\setminus v}x_{i}.$
@@
Let $Sp$ be an unbalanced split graph. Then, we have the following:
@@
    It is the content of Theorem 4.4.
@@
    It follows from Theorem 4.9 and Section 4(1).
@@
    From Section 4(2) we easily derive that the intersection of all the maximum independent sets of $Sp$ is $S^{*}$. Hence, $\operatorname{Supp}(Sp)\subset S^{*}$ by Theorem 4.9, and so $W\cap\operatorname{Supp}(Sp)=\varnothing$ by (1).
    Alternatively: if $x\in\operatorname{\mathcal{N}}(Sp)$, then $x_{u}=x_{v}$ for all $u,v\in W$, by Section 4 and Theorem 4.4. Since $\operatorname{Supp}(Sp)$ is independent by Theorem 4.9, it must be $x_{w}=0$ for all $w\in W$, because $|W|\geq 2$ and $W$ is a clique. Consequently, $W\cap\operatorname{Supp}(Sp)=\varnothing$.
  5. (5)
    If $u,v$ are distinct swing vertices, observe that $A(Sp)_{*u}=A(Sp)_{*v}$. As a consequence, $e_{u}-e_{v}\in\operatorname{\mathcal{N}}(Sp)$. Then, $W\subset\operatorname{Supp}(Sp)$.
    On the other hand, we know from Section 4(3) that $S^{*}\cup W$ is the unique maximum independent set of $Sp$. Therefore, $\operatorname{Supp}(Sp)\subset S^{*}\cup W$, by Theorem 4.9. ∎
@@ 5 Split graphs with supported clique vertices
In this section we prove that the subspace of kernel vectors with support in the clique (the clique-kernel) is at most one-dimensional. This result leads to a precise dichotomy: the overall nullity of the split graph is either equal to the nullity of its biadjacency matrix $R$, or exactly one greater. This establishes a complete and computable classification of the nullity based purely on the dimension of this key subspace, tightly linking the combinatorial presence of supported clique vertices to the algebraic property of the graph’s nullity.
The following result is crucial as it shows that when clique vertices appear in the support, they do so in essentially a one-dimensional way.
@@
Now we can show that the appearance of clique vertices in the support is an ”all-or-nothing” phenomenon in a one-dimensional sense.
@@
We obtain the desired result by combining item (3) of Theorem 3.4 with Section 5. ∎
@@
By Section 5, $\dim(\operatorname{Cker}{\left(Sp\right)})\leq 1$. If $\dim(\operatorname{Cker}{\left(Sp\right)})=0$, then by Theorem 3.4(3), $\operatorname{Supp}(Sp)\subset S$, and hence $\operatorname{nl}(Sp)=\operatorname{nl}(R)$ by Theorem 3.2(3). If $\dim(\operatorname{Cker}{\left(Sp\right)})=1$, then by Theorem 5.2, $\operatorname{Supp}(Sp)\cap K\neq\varnothing$, and hence $\operatorname{nl}(Sp)=\operatorname{nl}(R)+1$ by Theorem 5.3. ∎
@@ 6 Split graphs with nullity 1
Having determined the general structure of the nullspace, we now specialize in split graphs whose nullity is exactly one. This condition imposes a remarkable rigidity on the graph’s structure, forcing a highly constrained relationship between the kernel, the adjugate of the adjacency matrix, and the graph’s support.
In this section, we leverage the classical properties of the adjugate for singular matrices to derive precise combinatorial descriptions of these graphs, culminating in formulas that link the coordinates of the unique kernel vector to the determinants of principal submatrices.
@@
So, $a_{i}\in\operatorname{\mathcal{N}}(A)$ for all $i\in[n]$, from which we deduce that $\operatorname{rk}(\operatorname{adj}(A))\leq 1$, because $\operatorname{nl}(A)=1$. On the other hand, we know by Section 6 that $\operatorname{rk}(\operatorname{adj}(A))\neq 0$. Therefore, $\operatorname{rk}(\operatorname{adj}(A))=1$. ∎
@@
    Let $\operatorname{\mathcal{N}}(G)=\langle z\rangle$ and $s=|\operatorname{Supp}(G)|$. First, we label $V(G)$ using the ordering $(\operatorname{Supp}(G),V(G)\setminus\operatorname{Supp}(G))$. By this labeling, $z=(x,0)^{t}$, where $x\in\mathbb{R}^{s}$ and $x_{v}\neq 0$ for all $v\in[s]$. Then, by Section 6, we can write
@@
From Section 6, we know that $B_{i}$ spans the subspace $\langle\{A_{*j}:j\in\operatorname{Supp}(A)\}\rangle$. With this in mind, it is clear that $B$ spans $\operatorname{\mathcal{R}}(A)$. Since $|B|=n-1=\operatorname{rk}(A)$, we can finally conclude that $B$ is a basis for $\operatorname{\mathcal{R}}(A)$. ∎
@@
This theorem provides a very concrete and useful result: for any graph with nullity 1 and no isolated vertices, you can obtain a basis for the column space by simply removing any one column corresponding to a supported vertex.
@@
Since $G$ has no isolated vertices, $A(G)_{*j}\neq 0$ for all $j\in V(G)$. If $\operatorname{Supp}(G)=\{u\}$ for some $u$, then $\operatorname{\mathcal{N}}(G)=\langle e_{u}\rangle$. But, $0=A(G)e_{u}=A(G)_{*u}$, contradicting that all the columns of $A(G)$ are nonzero. Therefore, $|\operatorname{Supp}(G)|\geq 2$, and so we can apply Section 6. ∎
@@
Since $P$ is obtained by selecting a basis for $\operatorname{\mathcal{R}}(R)$, we have $\operatorname{\mathcal{R}}(P)=\operatorname{\mathcal{R}}(R)$ (hence, $\operatorname{\mathcal{R}}((I-J)^{-1}P)=\operatorname{\mathcal{R}}((I-J)^{-1}R)$) and $\operatorname{\mathcal{N}}(P^{t})=\operatorname{\mathcal{N}}(R^{t})$. Therefore, $\operatorname{Cker}{\left(Sp\right)}=\operatorname{Cker}{\left(Sp^{\prime}\right)}$. Since $\operatorname{Cker}{\left(Sp\right)}\neq\{0\}$, by Theorem 3.4(3), we have $\operatorname{Supp}(Sp^{\prime})$ $\cap K\neq\varnothing$. Since $\operatorname{\mathcal{N}}(P)=\{0\}$, applying Section 6 to $Sp^{\prime}$ gives $\operatorname{nl}(Sp^{\prime})=1$. ∎
@@
    The equality $\operatorname{Supp}(Sp^{\prime})\cap K=\operatorname{Supp}(Sp)\cap K$ can easily be derived from Section 6. To obtain that $\operatorname{Supp}(Sp^{\prime})\cap S^{\prime}\subset\operatorname{Supp}(Sp)\cap S^{\prime}$, recall that, by item (2) of Theorem 3.4 and item (1) of Theorem 3.2, there exists a (unique) vector $y_{0}$ such that $(I-J)^{-1}Py_{0}=z$. Then,
@@
    Let $M$ be the $|V(Sp)|\times(|V(Sp^{\prime})|-1)$ matrix obtained from $A(Sp)$ by deleting its last $|S|-\operatorname{rk}(R)$ columns and the column $s$. Since $\operatorname{rk}(M)$ is the size of its largest invertible submatrix, we have that $\operatorname{rk}(M)=|V(Sp^{\prime})|-1$ because $M$ contains $A(Sp^{\prime}-s)$, which is nonsingular by Section 6 and Theorem 6.3. Hence, $B_{s}$ is a linearly independent set.
@@ 7 Tyshkevich composition
Understanding how the nullspace behaves under graph operations is a natural and important question. In this section, we analyze how the nullspace of a split graph interacts with the Tyshkevich composition, a well-known product that builds a larger graph by fully joining the clique of a split graph to an arbitrary second graph.
@@
This operation is relevant for studying the nullspace of split graphs, as it allows us to build more complex split graphs from simpler components while potentially controlling spectral properties.
@@ 8 Determinant of split graphs
Thanks to the block structure of their adjacency matrices, split graphs admit explicit combinatorial determinant formulas that can be obtained with the Schur complement and Cauchy’s formula.
@@
Theorem 8.2 (Cauchy’s formula, [10]).
If $M$ is a square matrix and $u,v$ are vectors, then
$\det(M+uv^{t})=\det(M)+v^{t}\operatorname{adj}(M)u.$
@@
$\det(Sp)=(-1)^{k-1}\left((k-1)\det(R^{t}R)-r^{t}\operatorname{adj}(R^{t}R)r\right).$
@@
Using the well known Sherman–Morrison formula, $(I-J)^{-1}=I-(k-1)^{-1}J$, we have:
@@
Finally, we apply Theorem 8.2 with $M=R^{t}R,u=\frac{r}{1-k}$ and $v=r$:
$\det\left(R^{t}R-\frac{rr^{t}}{k-1}\right)=\det(R^{t}R)-\frac{r^{t}\operatorname{adj}(R^{t}R)r}{k-1}.\qed$
Notice that the vector $r=R^{t}\operatorname{1\!\!1}$ in Theorem 8.3 is the degree sequence of the independent part of $Sp(K,S)$, i.e., $r_{v}=\deg_{Sp}(v)$, for each $v\in S$.
@@
Let $Sp(K,S)$ be a split graph such that $k=|K|\geq 2$. If $r=R^{t}\operatorname{1\!\!1}$, then $Sp$ is singular if and only if
$(k-1)\det(R^{t}R)=r^{t}\operatorname{adj}(R^{t}R)r.$
In particular:
  1. 1.
    if $\operatorname{\mathcal{N}}(R)=\{0\}$, then $\det(Sp)=0$ if and only if $r^{t}(R^{t}R)^{-1}r=k-1$;
  2. 2.
    if $\operatorname{\mathcal{N}}(R)=\langle x\rangle$ for some $x\neq 0$, then $\det(Sp)=0$ if and only if $r^{t}x=0$;
  3. 3.
    if $\operatorname{nl}(R)\geq 2$, then $\det(Sp)=0$.
@@
The singularity characterization immediately follows from Theorem 8.3.
  1. (1).
    Since $R^{t}R$ is invertible, we have $\operatorname{adj}(R^{t}R)=(R^{t}R)^{-1}\det(R^{t}R)$.
  2. (2).
    Since $\operatorname{\mathcal{N}}(R^{t}R)=\operatorname{\mathcal{N}}(R)$, we have $\det(R^{t}R)=0$ and we can write $\operatorname{adj}(R^{t}R)=cxx^{t}$ for some $c\in\mathbb{R}\setminus 0$, by Section 6 and by the proof of Theorem 6.3. Therefore, $(k-1)0=r^{t}(cxx^{t})r=c(r^{t}x)(r^{t}x)^{t}$.
  3. (3).
    Since $\operatorname{nl}(R^{t}R)=\operatorname{nl}(R)\geq 2$, we have $\operatorname{adj}(R^{t}R)=0$ by Section 6 and $\det(R^{t}R)=0$. ∎
@@ Acknowledgments
The authors would like to thank the anonymous referee for their valuable comments and suggestions, which significantly improved Sections 4 and 8.